# Fundamental limits of weak learnability in high-dimensional multi-index models

### v1

@@ Abstract
Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural networks. This paper examines the theoretical boundaries of learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$. Our findings unfold in three parts: (i) first, we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!>\!0$; (ii) second, in the case where the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace consisting of directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$. The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$. In a limited but interesting set of really hard directions —akin to the parity problem— $\alpha_{c}$ is found to diverge. Finally, (iii) we demonstrate that interactions between different directions can result in an intricate hierarchical learning phenomenon, where some directions can be learned sequentially when coupled to easier ones. Our analytical approach is built on the optimality of approximate message-passing algorithms among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent.
@@ 1 Introduction
$y=g({{\bm{{W}}}}{\bm{x}}),\qquad{{\bm{{W}}}}\in\mathbb{R}^{p\times d},\quad k\in[p].$ (1)
@@
They define a rich class of hypotheses (Li, 1991), containing many widely studied functions in the statistical learning and theoretical computer science literature. The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z)\!=\!z\!+\!\sqrt{\Delta}Z$, with $Z\!\sim\!{\cal N}(0,1)$. In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, the terminology single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is employed when the link function is known and only the weights are learned. In this work, we use both interchangeability., $g(z)$ can be an arbitrary, possibility stochastic, link function, including some widely studied problems such as phase retrieval $g(z)\!=\!z^{2}$ and the perceptron $g(z)={\rm sign}(z)$. Some example of multi-index models include the Two-layer neural network ($p>1$), given by a link function $g(z_{1},\dots,z_{p})=\sum_{k=1}^{p}a_{k}\sigma(z_{k})$; Polynomial functions ($p>1$), given by any linear combination of products of $z_{1},\dots,z_{p}$, for example $g(z_{1},z_{2},z_{3})=1+z_{1}^{2}z_{2}^{2}-2z_{1}+z_{2}z_{3}^{3}$; and s-sparse parities ($p>1$) that can be embedded in a multi-index model by taking a link function $g(z_{1},\dots,z_{p})=\prod_{k\in I}{\rm sign}(z_{k})$ for any subset $I\subset[p]$ of size $|I|=s\leq p$. Training a multi-index model typically translates to a non-convex optimization problem. Several authors have thus used multi-index models as a test-bed for understanding the behavior of neural nets and gradient-descent in non-convex, high-dimensional contexts, e.g. (Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2024).
@@
To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models. This translates to the question of how many observations from model Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability. This can be studied both statistically (within the class of all, including exponential, algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms). In the single-index case ($p=1$), weak learnability has been heavily studied under probabilistic assumptions for the weights and data distribution (e.g. i.i.d. Gaussian or uniformly in the sphere). The statistical threshold for learnability has been characterised by Barbier et al. (2019) when the covariate dimension $d$ is large. Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020a) in the same regime. Similar results were also proven for other computational models: Damian et al. (2022) provided a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. Damian et al. (2024b) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
@@
  • We analyse the computational limits of weak learnability (as defined in Def. 2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\to\infty$ at fixed ratio $\alpha=\nicefrac{{n}}{{d}}$. Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural networks trained with gradient descent.
  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples in multi-index models: trivial directions can be learned with a single AMP iteration for any $\alpha>0$. When no such direction exists, we show there might be easy directions requiring $O(\log d)$ iterations and at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition.  [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold. In no way do we imply the existence or absence of a computational-to-statistical gap. We provide a sharp formula for the critical sample complexity $\alpha_{c}$ above which easy directions are weakly learnable. No efficient iterative algorithm can succeed for $\alpha<\alpha_{c}$. The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!O(d)$ samples.
  • We discuss how interactions among directions can lead to hierarchical learning phenomena. Interestingly, even hard directions can be learned when coupled to easier ones through a hierarchical process, which is reminiscent of the staircase phenomenon for one-pass SGD (Abbe et al., 2023). However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger. We fully characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
@@
The case $p=1$ has seen a lot of interest recently. In terms of gradient descent, Arous et al. (2021) has shown that if the link function $g$ is known, one-pass SGD achieves weak recovery in $n=\Theta(d^{\ell-1})$ steps, where $\ell$, known as the information exponent, is the first non-zero Hermite coefficient of the link function $g$. In the non-parametric setting —where $g$ is unknown— Berthier et al. (2024) has shown that a large-width two-layer neural network trained under one-pass SGD can learn an ”easy” single-index target ($\ell=1$) in $n=\Theta(d)$ steps. This is to be contrasted with full-batch GD, which can achieve weak recovery in $\Theta(1)$ steps with sample complexity $n=\Theta(d)$ even for particular problems with $\ell>1$ (Dandi et al., 2024b) (which is, indeed, the statistical optimal rate (Barbier et al., 2019; Mondelli and Montanari, 2019)). Similar results have also been proven for abstract computational models. Damian et al. (2022) has proven a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. More recently, Damian et al. (2024b) has proven a similar result for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$ and hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent. This turns out to be equivalent to the optimal computational weak recovery threshold of (Barbier et al., 2019; Mondelli and Montanari, 2019; Luo et al., 2019; Maillard et al., 2020b).
@@
While the situation is less understood, the multicase has also whitenessed a surge of recent interest. In particular, it has been used to understand the behavior of gradiend descent algorithm in neural networks. For instance, Abbe et al. (2022; 2023) showed that a certain class of staircase functions can be learned by large-width two-layer networks trained under one-pass SGD with sample complexity $n=\Theta(d)$. This is in stark contrast to (embedded) $s$-sparse parities, which require $n\geq d^{s-1}$ samples (Blum et al., 2003). Abbe et al. (2022; 2023) introduced the leap exponent, a direction-wise generalisation of the information exponent for multi-index models, and studied the class of staircase functions which can be efficiently learned with one-pass SGD. Bietti et al. (2023) showed that under a particular gradient flow scheme preserving weight orthogonality (a.k.a. Stiefel gradient flow), training follow a saddle-to-saddle dynamics, with the characteristic time required to escape a saddle given by the leap exponent. Interestingly, it was then shown that the limit discovered in these set of works could be bypassed by slightly different algorithm, for instance by smoothing the landscapes Damian et al. (2024a) or reusing batches multi-time Dandi et al. (2024b). The latter paper, in particular, showing that gradient descent could learn efficiently a larger class of multi-index models that previously believed to be possible. These findings highlight the need of a strict understanding of the limit learnability of these models.
@@
While these algorithm were studied in great detail for the single index models, see e.g. Barbier et al. (2019); Aubin et al. (2020b; a), and are at the roots of the spectral method that underline the learnability phase transition in this case Mondelli and Montanari (2019); Maillard et al. (2020b), much less is known in the multi-index case, with the exception of Aubin et al. (2019), who showed that particular instances of multi-index models known as committee machines can be learned with $n=\Theta(d)$ samples.
@@
We denote by $\mathcal{S}^{+}_{p}$, the cone of positive-semi-definite matrices in $\mathbb{R}^{p}$ and by $\succ$ the associated partial ordering. For ${{\bm{M}}}\in\mathcal{S}^{+}_{p}$, we denote by $\sqrt{M}$ the matrix-square root. |||||$\cdot$||| | ⋅ | | denotes the operator norm and $||\cdot||_{F}$ the Frobenius norm.
@@ 2 Settings and definitions
Given a covariate ${\bm{x}}\sim\mathcal{N}(0,\nicefrac{{1}}{{d}}{\bm{I}}_{d})$, we define the class of Gaussian multi-index models as likelihoods of the type:
$\mathbb{E}[y|{\bm{x}}]=g({{\bm{{W}}}}^{\star}{\bm{x}})$ (2)
where $g:\mathbb{R}^{p}\to\mathbb{R}$ is the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d. rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$. Note that this uniquely defines a joint distribution $p({\bm{x}},y)$ over $\mathbb{R}^{d+1}$.
@@
Let $V^{\star}\subset\mathbb{R}^{p}$ denote a subspace spanned by vectors representing components along ${{\bm{{W}}}}^{\star}$ such that each $\mathbf{v}\in V^{\star}$ maps to a vector $\mathbf{v}_{d}$ in $\mathbb{R}^{d}$ through the map $\mathbf{v}_{d}=({{\bm{{W}}}}^{\star})^{\top}\mathbf{v}$. Given an estimator $\hat{{{\bm{{W}}}}}\in\mathbb{R}^{p\times d}$ of ${{\bm{{W}}}}^{\star}$, we have weak recovery of a $V^{\star}$ if:
$\underset{{\bm{v}}\in V^{\star},\norm{{\bm{v}}}=1}{\inf}\norm{\frac{\hat{{{\bm{{W}}}}}({{\bm{{W}}}}^{\star})^{\top}{\bm{v}}}{d}}=\Theta_{d}(1).$ (3)
@@
Our main tool for characterising the computational bottlenecks in the Gaussian multi-index problem is an Approximate Message Passing (AMP) Algorithm 1 tailored to our problem. Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$. Therefore, it belongs to the class of first-order algorithms with linear running time on $O(nd+p^{3})$. The key property of AMP is that for a well-chosen ${\bm{g}}_{\rm out}$, it is provably optimal within the class of first-order methods (Celentano et al., 2020). This means that establishing learnability for the optimal AMP Algorithm 1 implies a computational lower bound in the class of first-order methods, which includes popular machine learning algorithms such as gradient descent (SG). For the Gaussian multi-index estimation problem 1, the optimal ${\bm{g}}_{\rm out}$ is simply given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
@@
where the conditional expectation is defined through the output channel $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$. Here ${\bm{\omega}}$ can be interpreted as an estimate of the pre-activations. The above denoiser can also be derived as an exact high-dimensional approximation of the Belief Propagation (BP) algorithm for the estimation of the marginals of the Bayesian posterior distribution:
@@
Our analysis is based on two remarkable properties that make AMP a particularly useful tool for studying high-dimensional estimation. The first property is that for any $t<O(\log{d})$ ), in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
@@
For multi-index models, the state evolution equations were derived by Aubin et al. (2020b) and rigorously proven by Gerbelot and Berthier (2023):
@@
Let $({\bm{x}}_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Consider running Algorithm 1 from initial condition $\hat{{{\bm{{W}}}}}^{0}\in\mathbb{R}^{p\times d}$, such that the initial overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ converges in probability to a limit ${{\bm{M}}}^{0}$ as $d\rightarrow\infty$. Denote by $\hat{{{\bm{{W}}}}}^{t}$ the resulting estimator at time $0\leq t\leq T$. Then, in the high-dimensional limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$ and $T<O(\log{d})$ ), the asymptotic mean-squared error on the label prediction is given by:
$\lim\limits_{n,d\to\infty}\mathbb{E}[(y-g(\hat{{{\bm{{W}}}}}^{t}{\bm{x}}))^{2}]=\mathbb{E}[(Y^{t}-g({\bm{Z}}))^{2}],$ $0\leq t\leq T$ (6)
where the expectation is taken over the following effective estimation process:
$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{\xi}}\right),\qquad{\bm{Z}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ (7)
with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$, and ${{\bm{M}}}^{t}$ is given by iterating the following state evolution equations from initial condition ${{\bm{M}}}^{0}$:

### v2

# Fundamental computational limits of weak learnability in high-dimensional multi-index models

@@ Abstract
Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural networks. This paper examines the theoretical boundaries of efficient learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$. Our findings unfold in three parts: (i) first, we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!>\!0$; (ii) second, in the case where the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace consisting of directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$. The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that it is conjectured that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$. In a limited but interesting set of really hard directions —akin to the parity problem— $\alpha_{c}$ is found to diverge. Finally, (iii) we demonstrate that interactions between different directions can result in an intricate hierarchical learning phenomenon, where some directions can be learned sequentially when coupled to easier ones. Our analytical approach is built on the optimality of approximate message-passing algorithms among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent.
@@ 1 Introduction
$y=g({{\bm{{W}}}}{\bm{x}}),\qquad{{\bm{{W}}}}\in\mathbb{R}^{p\times d},$ (1)
@@
where the mapping $g:\mathbb{R}^{p}\rightarrow\mathbb{R}$ can further be stochastic i.e. depend on independent noise $\xi\in\mathbb{R}^{p^{\prime}}$ for some finte $p^{\prime}$. They define a rich class of hypotheses (Li, 1991), containing many widely studied functions in the statistical learning and theoretical computer science literature. The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z,\xi)\!=\!z\!+\!\sqrt{\Delta}\xi$, with $\xi\!\sim\!{\cal N}(0,1)$. In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, the terminology single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is employed when the link function is known and only the weights are learned. In this work, we use both interchangeability., $g(z)$ can be an arbitrary, possibility stochastic, link function, including some widely studied problems such as phase retrieval $g(z)\!=\!z^{2}$ and the perceptron $g(z)={\rm sign}(z)$. Some example of multi-index models include the Two-layer neural network ($p>1$), given by a link function $g(z_{1},\dots,z_{p})=\sum_{k=1}^{p}a_{k}\sigma(z_{k})$; Polynomial functions ($p>1$), given by any linear combination of products of $z_{1},\dots,z_{p}$, for example $g(z_{1},z_{2},z_{3})=1+z_{1}^{2}z_{2}^{2}-2z_{1}+z_{2}z_{3}^{3}$; and s-sparse parities ($p>1$) that can be embedded in a multi-index model by taking a link function $g(z_{1},\dots,z_{p})=\prod_{k\in I}{\rm sign}(z_{k})$ for any subset $I\subset[p]$ of size $|I|=s\leq p$. Training a multi-index model typically translates to a non-convex optimization problem. Several authors have thus used multi-index models as a test-bed for understanding the behavior of neural nets and gradient-descent in non-convex, high-dimensional contexts, e.g. (Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2023).
@@
To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models. This translates to the question of how many observations from model Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability. This can be studied both statistically (within the class of all, including exponential time algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms). In the single-index case ($p=1$), weak learnability has been heavily studied under probabilistic assumptions for the weights and data distribution (e.g. i.i.d. Gaussian or uniformly in the sphere). The statistical threshold for learnability has been characterised by Barbier et al. (2019) when the covariate dimension $d$ is large. Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020) in the same regime. Similar results were also proven for other computational models: Damian et al. (2022) provided a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. Damian et al. (2024) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
@@
  • We analyse the computational limits of weak learnability (as defined in Def. 2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\to\infty$ at fixed ratio $\alpha=\nicefrac{{n}}{{d}}$ and constant number of indices $p=\Theta(1)$. Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural networks trained with gradient descent.
  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples in multi-index models: trivial directions can be learned with a single AMP iteration for any $\alpha>0$. When no such direction exists, we show there might be easy directions allowing recovery from arbitrarily small initial overlap $\epsilon>0$ in $\mathcal{O}(\log\frac{1}{\epsilon})$ steps with at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition. Heuristically, we expect recovery along easy directions from random initialization with $\alpha>\alpha_{c}$ in $\mathcal{O}(\log d)$ iterations. We further confirm this numerically through extensive simulations.  [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold. In no way do we imply the existence or absence of a computational-to-statistical gap. We provide a sharp formula for the critical sample complexity $\alpha_{c}$ above which easy directions are weakly learnable. It is conjectured that no efficient iterative algorithm can succeed for $\alpha<\alpha_{c}$. The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!\mathcal{O}(d)$ samples.
  • We discuss how interactions among directions can lead to hierarchical learning phenomena. Interestingly, even hard directions can be learned when coupled to easier ones through a hierarchical process, which is reminiscent of the staircase phenomenon for one-pass SGD (Abbe et al., 2023). However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger. We characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
@@
The case $p=1$ has seen a lot of interest recently. In terms of gradient descent, Arous et al. (2021) has shown that if the link function $g$ is known, one-pass SGD achieves weak recovery in $n=\Theta(d^{\ell-1})$ steps, where $\ell$, known as the information exponent, is the first non-zero Hermite coefficient of the link function $g$. In the non-parametric setting —where $g$ is unknown— Berthier et al. (2023) has shown that a large-width two-layer neural network trained under one-pass SGD can learn an ”easy” single-index target ($\ell=1$) in $n=\Theta(d)$ steps. This is to be contrasted with full-batch GD, which can achieve weak recovery in $\Theta(1)$ steps with sample complexity $n=\Theta(d)$ even for particular problems with $\ell>1$ (Dandi et al., 2024) (which is, indeed, the statistical optimal rate (Barbier et al., 2019; Mondelli and Montanari, 2019)). Similar results have also been proven for abstract computational models. Damian et al. (2022) has proven a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. More recently, Damian et al. (2024) has proven a similar result for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$ and hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent. This turns out to be equivalent to the optimal computational weak recovery threshold of (Barbier et al., 2019; Mondelli and Montanari, 2019; Luo et al., 2019; Maillard et al., 2020).
@@
While the situation is less understood, the multicase has also whitenessed a surge of recent interest. In particular, it has been used to understand the behavior of gradiend descent algorithm in neural networks. For instance, Abbe et al. (2022; 2023) showed that a certain class of staircase functions can be learned by large-width two-layer networks trained under one-pass SGD with sample complexity $n=\Theta(d)$. This is in stark contrast to (embedded) $s$-sparse parities, which require $n\geq d^{s-1}$ samples (Blum et al., 2003). Abbe et al. (2022; 2023) introduced the leap exponent, a direction-wise generalisation of the information exponent for multi-index models, and studied the class of staircase functions which can be efficiently learned with one-pass SGD. Bietti et al. (2023) showed that under a particular gradient flow scheme preserving weight orthogonality (a.k.a. Stiefel gradient flow), training follow a saddle-to-saddle dynamics, with the characteristic time required to escape a saddle given by the leap exponent. Interestingly, it was then shown that the limit discovered in these set of works could be bypassed by slightly different algorithm, for instance by smoothing the landscapes Damian et al. (2023) or reusing batches multi-time Dandi et al. (2024). The latter paper, in particular, showing that gradient descent could learn efficiently a larger class of multi-index models that previously believed to be possible. These findings highlight the need of a strict understanding of the limit learnability of these models.
@@
While these algorithm were studied in great detail for the single index models, see e.g. Barbier et al. (2019); Aubin et al. (2020b; a), and are at the roots of the spectral method that underline the learnability phase transition in this case Mondelli and Montanari (2019); Maillard et al. (2020), much less is known in the multi-index case, with the exception of Aubin et al. (2019), who showed that particular instances of multi-index models known as committee machines can be learned with $n=\mathcal{O}(d)$ samples.
@@
[…]
@@ 2 Settings and definitions
Given a covariate ${\bm{x}}\sim\mathcal{N}(0,\nicefrac{{1}}{{d}}{\bm{I}}_{d})$, we define the class of Gaussian multi-index models as output mappings of the type:
$y=g({{\bm{{W}}}}^{\star}{\bm{x}})$ (2)
where $g:\mathbb{R}^{p}\to\mathbb{R}$ denotes the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d. rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$. Note we allow the map $g:\mathbb{R}^{p}\to\mathbb{R}$ to be stochastic i.e $g(\cdot)=\tilde{g}(\cdot,{\bm{\xi}})$ for random variable ${\bm{\xi}}\in\mathbb{R}^{p}$ independent of $\mathbf{x}$ having a fixed dimension-independent law such that $\mathbb{E}\left[\norm{\xi}_{2}^{2}\right]<\infty$. [^3] Technically, our results hold for ${\bm{\xi}}\in\mathbb{R}^{p^{\prime}}$ with $p^{\prime}=\mathcal{O}(1)$, for example the case of a degenerated noise. But for simplicity of exposition we focus on the $p^{\prime}=p$ case.
@@
Let $V^{\star}\subset\mathbb{R}^{p}$ denote a subspace spanned by vectors representing components along ${{\bm{{W}}}}^{\star}$ such that each $\mathbf{v}\in V^{\star}$ maps to a vector $\mathbf{v}_{d}$ in $\mathbb{R}^{d}$ through the map $\mathbf{v}_{d}=({{\bm{{W}}}}^{\star})^{\top}\mathbf{v}$. Given an estimator $\hat{{{\bm{{W}}}}}\in\mathbb{R}^{p\times d}$ of ${{\bm{{W}}}}^{\star}$ with $\norm{\hat{{{\bm{{W}}}}}_{i}}^{2}_{F}=\Theta(d)$, we have weak recovery of a $V^{\star}$ if:
$\underset{{\bm{v}}\in\mathbb{R}^{p},\norm{{\bm{v}}}=1}{\inf}\norm{\frac{\hat{{{\bm{{W}}}}}({{\bm{{W}}}}^{\star})^{\top}{\bm{v}}}{d}}=\Theta(1).$ (3)
@@
Our main tool for characterising the computational bottlenecks in the Gaussian multi-index problem is an Approximate Message Passing (AMP) Algorithm 1 tailored to our problem. Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$ (defined later in Equation 4). Therefore, it belongs to the class of first-order algorithms with linear running time of $\mathcal{O}(nd)$. Intuitively, AMP estimates the mean ${\bm{\omega}}$ and variance ${\bm{V}}$ of ${{\bm{{W}}}}^{*}{\bm{x}}$ and the mean $\hat{{{\bm{{W}}}}}$ and covariance $\hat{{\bm{C}}}$ of ${{\bm{{W}}}}^{*}$ through a fixed point iteration described in Algorithm 1. The key property of AMP is that for a well-chosen ${\bm{g}}_{\rm out}$, it is provably optimal within the class of first-order methods (Celentano et al., 2020). This means that establishing learnability for the optimal AMP Algorithm 1 implies a computational lower bound in the class of first-order methods, which includes popular machine learning algorithms such as gradient descent (SG). For the Gaussian multi-index estimation problem 1, the optimal ${\bm{g}}_{\rm out}$ is simply given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
@@
where the conditional expectation is defined through the output channel $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$. For stochastic mappings $g(\cdot)=g(\cdot,\xi)$, the expectation is w.r.t both $Z,\xi$ conditioned on $Y$. Here ${\bm{\omega}}$ can be interpreted as an estimate of the pre-activations.
Our results rely on the following assumption on the above denoiser, translating to an indirect assumption on $g,\xi$:
Assumption 1.
${\bm{g}}_{\rm out}:\mathbb{R}^{p^{2}+p+1}\rightarrow\mathbb{R}^{p}\in\mathcal{C}^{2}$. Furthermore, the mapping $\hat{g}_{\rm out}:\mathbb{R}^{p+{p^{\prime}}}\rightarrow\mathbb{R}^{p}$ defined by ${\hat{g}}_{\rm out}(\mathbf{h},\xi)={\bm{g}}_{\rm out}(\tilde{g}((\mathbf{h},\xi),{\bm{\omega}},{\bm{V}})$ is pseudo-Lipschitz of finite-order for all ${\bm{\omega}},{\bm{V}}\in\mathbb{R}^{p},\mathbb{R}^{p\times p}$. Here, $\xi\in\mathbb{R}^{p^{\prime}}$ denotes independent noise satisfying $\mathbb{E}\left[\norm{\xi}^{2}\right]<\infty$ as per Definition 1.
The above denoiser can also be derived as an exact high-dimensional approximation of the Belief Propagation (BP) algorithm for the estimation of the marginals of the Bayesian posterior distribution:
@@
Our analysis is based on two remarkable properties that make AMP a particularly useful tool for studying high-dimensional estimation. The first property is that for any $t=\Theta(1)$, in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
@@
For multi-index models, the state evolution equations were derived by Aubin et al. (2020b) and rigorously proven by Gerbelot and Berthier (2023). The state evolution equations, in particular, provide an exact characterization of the asymptotic overlaps and prediction error.
@@
Let $({\bm{x}}_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Consider running Algorithm 1 from initial condition $\hat{{{\bm{{W}}}}}^{0}\in\mathbb{R}^{p\times d}$, such that the initial overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ converges in probability to a limit ${{\bm{M}}}^{0}$ as $d\rightarrow\infty$. Denote by $\hat{{{\bm{{W}}}}}^{t}$ the resulting estimator at time $0\leq t\leq T$. Then, in the high-dimensional limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, constant $p$ and any finite time $T$, the limiting overlaps satisfy:
[…]
with ${{\bm{M}}}^{t}$ satisfying the following state evolution equations from initial condition ${{\bm{M}}}^{0}$: